OpenStudy (anonymous):

Which set of line segments with the given lengths form a right triangle? 9, 27, 81 10, 24, 26 13, 15, 30 15, 18, 23

OpenStudy (cwrw238):

for right angled triangle c ^2 = a^2 + b^2 where c is the hypotenuse ( longest side)

OpenStudy (cwrw238):

so calculate each option eg 81^2 = 6561 27^2 = 729 9^2 =- 81 the last 2 results do not add up to 6561 so these line segments do not form right angled triangle

OpenStudy (anonymous):

I think it is 10, 24, 26

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw